OpenStudy (anonymous):

if then APR of an savings account is 7.2% and is compounded monthly what is the approximate APY of the account

OpenStudy (anonymous):

APY=(1+p)^(n)-1

OpenStudy (anonymous):

\[\left(1+\frac{.072}{12}\right)^{12}-1\]

OpenStudy (anonymous):

p= 0.072/12 n=1

OpenStudy (anonymous):

not getting the right answer the choices are A.1.07% B.7.44% C.10.74% D.6%

OpenStudy (anonymous):

will it be 7.44?

OpenStudy (anonymous):

seems reasonable http://www.wolframalpha.com/input/?i=%281%2B.072%2F12%29^%2812%29-1